OpenStudy (anonymous):

@jzellem1004 Okay here posted I will post the question so you dont forget (;

OpenStudy (anonymous):

mk ;]

OpenStudy (anonymous):

what are the numbers u want to use?

OpenStudy (anonymous):

2 and 4? lol idk they are my fave numbers

OpenStudy (anonymous):

@johnweldon1993 This is what I a currently working on (: I have to write an educational fairy tale lol I got the writing down I just need to answer the math questions to put in the fairy tale

OpenStudy (anonymous):

ok so I am not going to cheat it, so how would u graph tht

OpenStudy (anonymous):

Hm well i have to make a non-linear equation first and then I think I can graph it

OpenStudy (anonymous):

whoa i think i have to make more than 1

OpenStudy (anonymous):

no cuz y=m(x)+b

OpenStudy (anonymous):

Give me a minute to catch up here cause i am kinda confused

OpenStudy (anonymous):

iight ;]

OpenStudy (anonymous):

My stupid lesson has Linear equation yet my project is asking for non-linear >.<

OpenStudy (anonymous):

that's how the economy works now adays :] but don't worry I am here :P

OpenStudy (anonymous):

do you do cyber through riverside?

OpenStudy (anonymous):

no i dont know what that is

OpenStudy (anonymous):

nvm, it is a good thing ur hot

OpenStudy (anonymous):

did you just imply that I am dumb?

OpenStudy (anonymous):

no I am just saying lol

OpenStudy (anonymous):

I do virtual school but I live in FL so mine is different from whatever you guys have thats probably why idk what you are talking about

OpenStudy (anonymous):

o mk

OpenStudy (anonymous):

may I have a medal ;]

OpenStudy (johnweldon1993):

So let me give you a system of non-linear equations... \[\large y = x^2\] \[\large y = -x^2 + 8\] To solve this graphically....we graph the 2 functions... 

OpenStudy (johnweldon1993):